OpenStudy (anonymous):

A certain strain of virus grows in numbers at the rate of 50% per hour. If its present population is 88,000 what will be its population count in 6 hours?

OpenStudy (anonymous):

N = A(r)^(t/d) N = 88000(1.5)^(6/2)

OpenStudy (anonymous):

I think, I might be wrong.

OpenStudy (anonymous):

Follow along and try to notice the pattern The amount after 1 hour, A1 = 88000 + 88000*.5 = 88000(1+.5) The amount after 2 hours, A2 = A1 + A1*.5 = A1(1+.5) = 88000(1+.5)(1+.5) 3 hours, A3 = A2 + A2*.5 = A2(1+.5) = 88000(1+.5)(1+.5)(1+.5) 4 hours, A4 = A3 +A3*.5 = A3(1+.5) = 88000(1+.5)(1+.5)(1+.5)(1+.5) So to get An, I just start with 88000 and multiply by (1.5) n times, this is the same as multiplying by (1.5)^n

OpenStudy (anonymous):

Petewe, you are almost right. It's just that the power is 6, not 6/2.

OpenStudy (anonymous):

So what is the count?

OpenStudy (anonymous):

=(

OpenStudy (anonymous):

Hours | Population |Can also be written as 0 88000 88000 1 88000*(1.5) 88000*(1.5)^1 2 88000*(1.5)(1.5) 88000*(1.5)^2 3 88000*(1.5)(1.5)(1.5) 88000*(1.5)^3 4 88000*(1.5)(1.5)(1.5)(1.5) 88000*(1.5)^4 ... please notice the pattern and figure out how to find the population after 6 hours on your own